OpenStudy (anonymous):

Ho many three digit numbers have exactly two repeated digits? I found 270 is it right? let x be the repeated number we have three ways xx_ 9*10 numbers= 90 _xx 9*10=90 x_x 9*10 = 90 so 90*3= 270 three digit numbers . Is this right?

OpenStudy (reemii):

those numbers start at 100 and end at 999. (first digit can't be 0 !) other way of counting: for each starting number D, you can either repeat it or choose another digit that will be 'double'. D-D-E or D-E-E, or D-E-D. Therefore, "For each of the 9 possible choices for D, there are 3×9 ways to continue the number." I get 9×9×3.

OpenStudy (reemii):

243

OpenStudy (reemii):

the 27 other numbers are the one starting wth 0. you just overlooked this.

OpenStudy (anonymous):

No I think I considered that , so for example for XX_ the repeated x can be from 1 to 9 and the last number can be from 0 to 10 therefore 9*10 =90 3 digit number that look like 220,221,771 etc. then for _XX (e.g 577) the first digit can be from 1 to 9 and the doubled digit can be from 0 to 10 . so 9*10 = 90 . 100 , 111,122........900,911 the same for X_X , X can be from 1 to 9 and the middle can be 0 to 9 so we have 9*10 =90

OpenStudy (anonymous):

@reemii

OpenStudy (reemii):

you're not allowed to choose a _ digit that is already chosen as X. if you choose X=2, you can't choose 2 for the _, so you have only 9 choices.

OpenStudy (anonymous):

Oh my god , yes you are right. Thanks !!

OpenStudy (reemii):

np